OpenStudy (anonymous):

One last question! A line passes through the origin and has a slope of . Which of the following points does the line pass through? (0, ) (, 1) (1, 2) (2, 1)

OpenStudy (anonymous):

slope of?

OpenStudy (chosenmatt):

um your question...

OpenStudy (anonymous):

1/2, sorry about that CX

OpenStudy (anonymous):

I didn't see that until now, whoops.

OpenStudy (bibby):

Use point slope form \(y-y_1=m(x-x_1)\) where \((x_1,y_1)\)=(0,0) (the origin) and m is the slope

OpenStudy (anonymous):

so start at the origin

OpenStudy (anonymous):

Okay...

OpenStudy (bibby):

y-0=1/2(x-0) \(y=\frac{1}{2}x\) now plug in the respective x's and check if the y's match up or logically know that at x=0, y=0 x=1, y=1/2 x=2, y=1 etc.

OpenStudy (anonymous):

or the easiest way is to know that as y increases by 1 x increases by 2 going through the origin (0,0) so from that we can conclude that the next increment will be (2,1)

OpenStudy (anonymous):

Oh. Okay, I think I got it. Thanks y'all c: